# Hutchinson Trace Estimation for High-Dimensional and High-Order Physics-Informed Neural Networks

### 22 Dec 2023 v1

@@ 4.1 Nonlinear Second-Order PDEs with Anisotropic and Inseparable Solutions
Here are the implementation details. The model is a 4-layer fully connected network with 128 hidden units, which is trained via Adam [37] for 10K epochs in the two-body case and 20K epochs in the three-body case, respectively, with an initial learning rate 1e-3, which linearly decays to zero at the end of the optimization. We select 100 random residual points at each Adam epoch and 20K fixed testing points uniformly from the unit ball. We adopt the following model structure to satisfy the zero boundary condition with hard constraint and to avoid the boundary loss [38] $(1-\|\boldsymbol{x}\|_{2}^{2})u_{\theta}(\boldsymbol{x}),$ where $u_{\theta}(\boldsymbol{x})$ is the neural neural network model. We repeat our experiment five times with five independent random seeds and report the mean and standard deviation of the errors. For HTE, we choose $V=16$. For SDGD, we chose the batch size of dimension 16.
@@
Table 1: Computational results for Sine-Gordon equations with the two-body (Error_1) and three-body (Error_2) exact solutions.
The computational results for the two-body and three-body cases are presented in Table 1, where we compare vanilla PINN [45], SDGD [22], and HTE (Ours), in terms of speed (iteration per second), memory cost, and relative $L_{2}$ error for the two-body case (Error_1) and that for the three body case (Error_2). Note that when solving for the two PDEs with different solutions, the speed and the memory costs are similar, and only the error differs, so we only report speed and memory once.
@@ 4.2 High-Dimensional Gradient-Enhanced PINN
$L_{\text{PINN}}(\theta)=\frac{1}{2}\left(u_{\theta}(\boldsymbol{x})+\sin\left(u_{\theta}(\boldsymbol{x})\right)-g(\boldsymbol{x})\right)^{2}:=\frac{1}{2}r_{\theta}(\boldsymbol{x})^{2},$ (23)
@@
The implementation details are given as follows. The model is a 4-layer fully connected network with 128 hidden units, which is trained via Adam [37] for 10K epochs with an initial learning rate 1e-3, which linearly decays to zero at the end of the optimization. We select 100 random residual points at each Adam epoch and 20K fixed testing points uniformly from the unit ball. We adopt the following model structure to satisfy the zero boundary condition with hard constraint and to avoid the boundary loss [38] $(1-\|\boldsymbol{x}\|_{2}^{2})u_{\theta}(\boldsymbol{x}),$ where $u_{\theta}(\boldsymbol{x})$ is the neural neural network model. We repeat our experiment five times with five independent random seeds and report the mean and standard deviation of the errors. For HTE, we choose $V=16$ in all settings. We choose $\lambda_{\text{gPINN}}$ as ten in 100D and 1K D cases and choose it as ten thousand in 10K D and 100K D cases. The gPINN loss weight $\lambda_{\text{gPINN}}$ can be chosen to make the scale of the PINN loss and the scale of the gPINN loss similar during initialization, i.e., choose $\lambda_{\text{gPINN}}$ such that ${r}_{\theta}(\boldsymbol{x})^{2}\approx\lambda_{\text{gPINN}}\left\|\nabla_{\boldsymbol{x}}{r}_{\theta}(\boldsymbol{x})\right\|^{2}$ and $\hat{r}_{\theta}(\boldsymbol{x})^{2}\approx\lambda_{\text{gPINN}}\left\|\nabla_{\boldsymbol{x}}\hat{r}_{\theta}(\boldsymbol{x})\right\|^{2}$ at initialization.
@@
HTE PINN (Ours) Speed 1792.03it/s 1566.19it/s 1066.10it/s 345.10it/s
@@
HTE gPINN (Ours) Speed 616.79it/s 461.81it/s 229.18it/s 153.35it/s
@@
The computational results are presented in Table 2, where three algorithms are compared: vanilla PINN [45], vanilla gPINN [55], HTE-based PINN, and HTE-based gPINN. In the preliminary, we observe that both PINN and gPINN encounter failure when exceeding the 80GB memory limit of the A100 GPU in dimensions surpassing 1000D. Conversely, PINN and gPINN based on HTE successfully operate in dimensions as high as 100K, demonstrating swift performance. Comparing PINN and gPINN, we note an enhancement in the latter’s performance improvement, albeit at a slightly slower convergence pace, due to the additional first-order derivatives of the residual. The advantages of gPINN become more pronounced in higher dimensions (exceeding 1000D), leading to an additional order of magnitude reduction in PINN error. However, in scenarios with 100D and 1000D, HTE-based gPINN performs slightly inferior to conventional gPINN. This arises because HTE is ultimately an approximation, and in high-dimensional gPINN, it introduces $d$ additional losses, where $d$ is the dimensionality. Hence, the approximation with HTE tends to incur more significant errors. Nevertheless, the performance of HTE-based gPINN remains comparable to conventional gPINN. In summary, we verify that HTE can accelerate and scale up gPINN to very high dimensions. Moreover, gPINN continues to enhance the performance of PINN in high-dimensional PDEs. To the best of our knowledge, previous work [55] only demonstrated gPINN’s effectiveness in low-dimensional PDEs, while our work is the first to show that in high dimensions.
@@ 4.3 Biharmonic Equation
Table 3: Computational results for the Biharmonic equation.
@@ Appendix A Proof
Convergence of $L_{\text{HTE}}(\theta;\{\boldsymbol{v}_{i}\}_{i=1}^{V})$. Due to the strong law of large numbers,
@@
Unbiasedness of $L_{\text{HTE, unbiased}}(\theta;\{\boldsymbol{v}_{i},\hat{\boldsymbol{v}}_{i}\}_{i=1}^{V})$. Since HTE is an unbiased trace estimator,
@@
For the unit Gaussian distribution, we have that
$=3\Delta^{2}u(\boldsymbol{x}).$
A.3 Proof of Theorem 3.2
@@
where $\binom{d}{B}=\frac{d(d-1)\cdots(d-B+1)}{B(B-1)\cdots 1}$.
A.4 Proof of Theorem 3.3
@@
With the HTE estimator of batch size $V$, the variance is
$\mathbb{V}\left[\frac{1}{V}\sum_{k=1}^{V}\boldsymbol{v}_{k}^{\mathrm{T}}A\boldsymbol{v}_{k}\right]=\frac{1}{V}\sum_{i\neq j}A_{ij}^{2}.$ (36)

### 03 Mar 2024 v2

@@ 4.1 Nonlinear Second-Order PDEs with Anisotropic and Inseparable Solutions
Here are the implementation details. The model is a 4-layer fully connected network with 128 hidden units activated by Tanh, which is trained via Adam [38] for 10K epochs in the two-body case and 20K epochs in the three-body case, respectively, with an initial learning rate 1e-3, which linearly decays to zero at the end of the optimization. We select 100 random residual points at each Adam epoch and 20K fixed testing points uniformly from the unit ball. We adopt the following model structure to satisfy the zero boundary condition with hard constraint and to avoid the boundary loss [39] $(1-\|\boldsymbol{x}\|_{2}^{2})u_{\theta}(\boldsymbol{x}),$ where $u_{\theta}(\boldsymbol{x})$ is the neural neural network model. We repeat our experiment five times with five independent random seeds and report the mean and standard deviation of the errors. For HTE, we choose $V=16$. For SDGD, we chose the batch size of dimension 16.
@@
Table 1: Computational results for Sine-Gordon equations with the two-body (Error_1) and three-body (Error_2) exact solutions. Under the same computational constraints, SDGD [22] and our HTE perform similarly, and they are all competitive against vanilla PINNs [46].
The computational results for the two-body and three-body cases are presented in Table 1, where we compare vanilla PINN [46], SDGD [22], and HTE (Ours), in terms of speed (iteration per second), memory cost (MB), and relative $L_{2}$ error for the two-body case (Error_1) and that for the three body case (Error_2). Note that when solving for the two PDEs with different solutions, the speed and the memory costs are similar, and only the error differs, so we only report speed and memory once. We use MB as the memory unit because it is more precise. Typically, when checking the GPU memory using the Linux command “nvidia-smi”, the output is in MB, and converting it to GB may result in multiple decimal places, which is not as clear and accurate as a whole number in MB.
@@
4.1.1 Additional Study 1: Effect of $V$ on HTE’s Convergence
$V$ 1 5 10 15 16
Speed 435.63it/s 402.10it/s 382.91it/s 350.18it/s 345.10it/s
Memory 1073MB 1075MB 1080MB 1085MB 1089MB
Error_1 3.02E-03$\pm$2.16E-03 2.70E-03$\pm$1.35E-03 2.55E-03$\pm$1.08E-03 2.41E-03$\pm$8.10E-04 2.38E-3$\pm$1.72E-3
Error_2 2.23E-05$\pm$5.00E-05 1.96E-05$\pm$2.91E-05 1.75E-05$\pm$3.28E-05 1.62E-05$\pm$1.25E-05 1.59E-5$\pm$9.29E-6
Table 2: Effects of the number of random vectors $V$ on the convergence for 100,000-dimensional Sine-Gordon equations with the two-body (Error_1) and three-body (Error_2) exact solutions.
We conduct additional experiments for the case of the highest 100,000-dimensional Sine Gordon equation by following Section 4.1 in the main paper, except with different $V$. We aim to understand the effect of HTE batch size $V$ on its convergence, speed, and memory usage. Table 2 shows that $V=1$ is sufficient for a good convergence and that the error is generally reduced by increasing the HTE batch size $V$. Increasing the batch size results in slightly larger memory consumption and slower speed, as larger batch sizes entail greater computational load. These results demonstrate the robustness of HTE under minimal batch size $V$ and that we can always improve HTE by increasing $V$ at the cost of more memory usage and slower speed. This result also demonstrates the convergence rate of HTE towards that of the full PINN as $V$ increases, noting that we observe that HTE with $V=16$ is essentially indistinguishable from the PINN.
The philosophy is that optimizing and learning part of the dimensions is enough to achieve good performances in high-dimensional PDEs. In the PINN community, people are sampling residual points to accelerate convergence and reduce computational cost. HTE is, in essence, sampling dimensions while keeping a minibatch stochastic gradient.
The design philosophy of HTE is also akin to stochastic gradient descent (SGD) in machine learning. Given a large number of data points, we do not compute and optimize the loss for each data point at every iteration. Instead, we select a minibatch to construct an unbiased gradient, achieving good results. Today, SGD is the standard method in machine learning. Similarly, in high-dimensional PDEs, we do not need to optimize each dimension at every iteration. The randomized gradient obtained through HTE is sufficient.
4.1.2 Additional Study 2: Effect of Biased and Unbiased Versions of HTE
Method Metric 100 D 1,000 D 5, [^0] D 10,000 D 100,000 D
Biased HTE Speed 1792.03it/s 1566.19it/s 1275.81it/s 1066.10it/s 345.10it/s
Memory 869MB 890MB 913MB 927MB 1089MB
Unbiased HTE Speed 1619.81it/s 1437.20it/s 1139.63it/s 958.64it/s 303.73it/s
Memory 881MB 901MB 920MB 932MB 1094MB
Error_1 6.21E-3$\pm$3.01E-3 1.23E-3$\pm$4.02E-4 2.48E-3$\pm$1.50E-3 1.79E-3$\pm$8.24E-4 2.33E-3$\pm$1.67E-3
Error_2 7.43E-3$\pm$5.02E-4 8.72E-4$\pm$1.09E-5 1.85E-4$\pm$2.54E-6 9.84E-5$\pm$3.92E-6 1.45E-5$\pm$1.12E-5
Table 3: Effects of the biased and unbiased versions of HTE with $V=16$ on the convergence for Sine-Gordon equations with the two-body (Error_1) and three-body (Error_2) exact solutions.
Table 3 presents additional experimental results for comparing the biased and unbiased versions of HTE. We conducted experiments following the first Sine-Gordon equations presented in the main text, utilizing the same hyperparameters, except we compared the biased and unbiased versions. The results indicate that the unbiased version is approximately 10% slower than the biased version, with slightly higher memory consumption and a minor improvement in error. The increased cost and slower speed of the unbiased version stem from the necessity to compute HTE twice using two groups of independent $\{\boldsymbol{v}_{i}\}_{i=1}^{V}$ and $\{\hat{\boldsymbol{v}_{i}}\}_{i=1}^{V}$, whereas the biased version only requires one computation for $\{\boldsymbol{v}_{i}\}_{i=1}^{V}$. However, while the unbiased version’s unbiased gradients lead to a slight improvement in performance, in most cases, the biased version’s performance is already sufficiently good.
@@ 4.2 High-Dimensional Gradient-Enhanced PINN
$L_{\text{PINN}}(\theta)=\frac{1}{2}\left(\Delta u_{\theta}(\boldsymbol{x})+\sin\left(u_{\theta}(\boldsymbol{x})\right)-g(\boldsymbol{x})\right)^{2}:=\frac{1}{2}r_{\theta}(\boldsymbol{x})^{2},$ (23)
@@
The implementation details are given as follows. The model is a 4-layer fully connected network with 128 hidden units activated by Tanh, which is trained via Adam [38] for 10K epochs with an initial learning rate 1e-3, which linearly decays to zero at the end of the optimization. We select 100 random residual points at each Adam epoch and 20K fixed testing points uniformly from the unit ball. We adopt the following model structure to satisfy the zero boundary condition with hard constraint and to avoid the boundary loss [39] $(1-\|\boldsymbol{x}\|_{2}^{2})u_{\theta}(\boldsymbol{x}),$ where $u_{\theta}(\boldsymbol{x})$ is the neural neural network model. We repeat our experiment five times with five independent random seeds and report the mean and standard deviation of the errors. For HTE, we choose $V=16$ in all settings. We choose $\lambda_{\text{gPINN}}$ as ten in 100D and 1K D cases and choose it as ten thousand in 10K D and 100K D cases. The gPINN loss weight $\lambda_{\text{gPINN}}$ can be chosen to make the scale of the PINN loss and the scale of the gPINN loss similar during initialization, i.e., choose $\lambda_{\text{gPINN}}$ such that ${r}_{\theta}(\boldsymbol{x})^{2}\approx\lambda_{\text{gPINN}}\left\|\nabla_{\boldsymbol{x}}{r}_{\theta}(\boldsymbol{x})\right\|^{2}$ and $\hat{r}_{\theta}(\boldsymbol{x})^{2}\approx\lambda_{\text{gPINN}}\left\|\nabla_{\boldsymbol{x}}\hat{r}_{\theta}(\boldsymbol{x})\right\|^{2}$ at initialization.
@@
HTE PINN (Ours) Memory 869MB 890MB 927MB 1089MB
Speed 1792.03it/s 1566.19it/s 1066.10it/s 345.10it/s
@@
HTE gPINN (Ours) Memory 869MB 890MB 927MB 1089MB
Speed 616.79it/s 461.81it/s 229.18it/s 153.35it/s
@@
The computational results are presented in Table 4, where three algorithms are compared: vanilla PINN [46], vanilla gPINN [58], HTE-based PINN, and HTE-based gPINN. In the preliminary, we observe that both PINN and gPINN encounter failure when exceeding the 80GB memory limit of the A100 GPU in dimensions surpassing 1000D. Conversely, PINN and gPINN based on HTE successfully operate in dimensions as high as 100K, demonstrating swift performance. Comparing PINN and gPINN, we note an enhancement in the latter’s performance improvement, albeit at a slightly slower convergence pace, due to the additional first-order derivatives of the residual. We also observe the same memory cost for PINN and gPINN since gPINN necessitates an additional gradient concerning the input $\boldsymbol{x}$ for the PINN or HTE residual, and we accomplish this computation using forward-mode automatic differentiation. Forward mode is highly memory efficient, so it does not result in any additional memory consumption. Therefore, both PINN and gPINN have the same memory consumption. However, forward-mode computation does require more time. The advantages of gPINN become more pronounced in higher dimensions (exceeding 1000D), leading to an additional order of magnitude reduction in PINN error. However, in scenarios with 100D and 1000D, HTE-based gPINN performs slightly inferior to conventional gPINN. This arises because HTE is ultimately an approximation, and in high-dimensional gPINN, it introduces $d$ additional losses, where $d$ is the dimensionality. Hence, the approximation with HTE tends to incur more significant errors. Nevertheless, the performance of HTE-based gPINN remains comparable to conventional gPINN. In summary, we verify that HTE can accelerate and scale up gPINN to very high dimensions. Moreover, gPINN continues to enhance the performance of PINN in high-dimensional PDEs. To the best of our knowledge, previous work [58] only demonstrated gPINN’s effectiveness in low-dimensional PDEs, while our work is the first to show that in high dimensions.
@@ 4.3 Biharmonic Equation
Table 5: Computational results for the Biharmonic equation. In all-dimensional biharmonic PDEs, HTE consistently achieves results similar to those of full PINN while significantly outpacing full PINN in terms of speed and notably reducing memory consumption. In the case of 200 dimensions, where full PINN exceeds memory limits, HTE still converges rapidly.
@@ Appendix A Proof
Theorem A.1.
[…]
Proof.
Convergence of $L_{\text{HTE}}(\theta;\{\boldsymbol{v}_{i}\}_{i=1}^{V})$. Due to the strong law of large numbers and the fact that HTE is unbiased, i.e., $\mathbb{E}_{\{\boldsymbol{v}_{i}\}_{i=1}^{V}}\left[\frac{1}{V}\sum_{i=1}^{V}\boldsymbol{v}_{i}^{\mathrm{T}}A_{\theta}(\boldsymbol{x},t)\boldsymbol{v}_{i}\right]=\operatorname{Tr}\left(A_{\theta}(\boldsymbol{x},t)\right)$,
@@
Unbiasedness of $L_{\text{HTE, unbiased}}(\theta;\{\boldsymbol{v}_{i},\hat{\boldsymbol{v}}_{i}\}_{i=1}^{V})$. Since HTE is an unbiased trace estimator, i.e.,
$\quad\mathbb{E}_{\{\boldsymbol{v}_{i}\}_{i=1}^{V}}\left[\frac{1}{V}\sum_{i=1}^{V}\boldsymbol{v}_{i}^{\mathrm{T}}A_{\theta}(\boldsymbol{x},t)\boldsymbol{v}_{i}+B_{\theta}(\boldsymbol{x},t)\right]$ (33)
$=\mathbb{E}_{\{\hat{\boldsymbol{v}}_{i}\}_{i=1}^{V}}\left[\frac{1}{V}\sum_{i=1}^{V}\hat{\boldsymbol{v}}_{i}^{\mathrm{T}}A_{\theta}(\boldsymbol{x},t)\hat{\boldsymbol{v}}_{i}+B_{\theta}(\boldsymbol{x},t)\right]=\operatorname{Tr}\left(A_{\theta}(\boldsymbol{x},t)\right)+B_{\theta}(\boldsymbol{x},t).$ (34)
The random variables $\{\boldsymbol{v}_{i}\}_{i=1}^{V}$ and $\{\hat{\boldsymbol{v}}_{i}\}_{i=1}^{V}$ are also independent. Thus
@@
Theorem A.2.
(Theorem 3.2 in the main text) Given the SDGD estimator of the trace $\operatorname{Tr}(A)\approx\frac{d}{B}\sum_{i\in I}A_{ii}$ given an index set $I\subset\{1,2,\cdots,d\}$ whose cardinality $|I|=B$, where $B$ is the SDGD’s batch size for the dimension, its variance is
where $\binom{d}{B}=\frac{d(d-1)\cdots(d-B+1)}{B(B-1)\cdots 1}$.
Proof.
@@
where $\binom{d}{B}=\frac{d(d-1)\cdots(d-B+1)}{B(B-1)\cdots 1}$. ∎
A.3 Proof of Theorem 3.3
Theorem A.3.
(Theorem 3.3 in the main text) Given the HTE estimator of the trace $\operatorname{Tr}(A)\approx\frac{1}{V}\sum_{k=1}^{V}\boldsymbol{v}_{k}^{\mathrm{T}}A\boldsymbol{v}_{k},$ where $V$ is the HTE batch size and each dimension of $\boldsymbol{v}_{k}\in\mathbb{R}^{d}$ is an $i.i.d.$ sample from the Rademacher distribution, then its variance is $\frac{1}{V}\sum_{i\neq j}A_{ij}^{2}.$
Proof.
@@
With the HTE estimator of batch size $V$ and $\boldsymbol{v}_{k}$ are independent for different $k$, the variance is
$\mathbb{V}\left[\frac{1}{V}\sum_{k=1}^{V}\boldsymbol{v}_{k}^{\mathrm{T}}A\boldsymbol{v}_{k}\right]=\frac{1}{V^{2}}\sum_{k=1}^{V}\mathbb{V}\left[\boldsymbol{v}_{k}^{\mathrm{T}}A\boldsymbol{v}_{k}\right]=\frac{1}{V^{2}}\sum_{k=1}^{V}\left[\sum_{i\neq j}A_{ij}^{2}\right]=\frac{1}{V}\sum_{i\neq j}A_{ij}^{2}.$ (40)
A.4 Proof of Theorem 3.4
Theorem A.4.
(Theorem 3.4 in the main text) The biharmonic operator can be unbiasedly estimated using the following TVPs:
where $\mathcal{N}(0,I)$ is the $d$-dimensional unit Gaussian, and
and the TVP is defined as
Proof.
For the unit Gaussian distribution, we have that
$=3\Delta^{2}u(\boldsymbol{x}).$
Here, we use the fact that for the one-dimensional Gaussian random variable $v\sim\mathcal{N}(0,1)$, its second moment $\mathbb{E}_{v}[v^{2}]=1$ and its fourth moment $\mathbb{E}_{v}[v^{4}]=3$. Note that each entry of the vector $\boldsymbol{v}$ is the one-dimensional Gaussian. ∎